A 0.070 0-kg of brass is heated to 230.0°C and then dropped into a beaker containing 0.900 kg of water. If the final equilibrium temperature of the mixed system is 23.4°C, find the initial temperature of water.


1
Expert's answer
2021-05-05T07:45:23-0400

"(C_pm \triangle T)_{water}=(C_pm \triangle T)_{brass}"

"4186 \times 0.9 \times 195=920 \times 0.07 \times \triangle T"

"920\cdot \:0.07 \triangle T=4186\cdot \:0.9\cdot \:195"

"6440 \triangle T=73464300"

"\triangle T=11384.1"

"T_1 = 23.4-11407.5=-11384.1^0C"